A charge 4.98 nC is placed at the origin of an xy-coordinate system, and a charge -1.96 nC is placed on the positive x-axis at x = 3.97 cm . A third particle, of charge 6.00 nC is now placed at the point x = 3.97 cm , y = 3.00 cm .

Part A

Find the x-component of the total force exerted on the third charge by the other two.

Part B

Find the y-component of the total force exerted on the third charge by the other two.

Part C

Find the magnitude of the total force acting on the third charge.

Part D

Find the direction of the total force acting on the third charge.
Answer #1

force between q1 and q3

F13 = k q1 q3 / r13^2 = 9*10^9* 4.98* 6*10^-18 / (0.0397^2 + 0.03^2)

F13 = 1.086*10^-4 N

direction of F1 with horizontal

x = arctan ( 3/ 3.97) = 37.08

force between q2 and q3

F23 = 9* 10^9* 1.96*10^-18* 6 / 0.03^2

F23 = 1.176*10^-4 N

=============

a)

net horizontal component of force

Fx = F13 cos x = 8.66*10^-5 N

=============

b)

net vertical component of force

Fy = F13 sin x - F23

Fy = - 5.21 *10^-5 N

============

c)

net force acting

F^2 = Fx^2 + Fy^2

F = 10.1 *10^-5 N

============

d)

direction measured from +x in cw way

phi = arctan ( Fy/ Fx)

phi = arctan ( 5.21 / 8.66)

phi = 31

============

Comment in case any doubt, will reply for sure.. Goodluck